# Asymptotically Unbiased Synthetic Control Methods by Density Matching

### v4

@@ 5.1 Identification and Estimation Strategy
Section 4 and the results in Ferman and Pinto (2021) indicate that combining
@@
We begin by positing that the distribution of the target unit’s outcome $Y^{N}_{0,t}$ can be expressed as a linear combination of the distributions of $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. For simplicity, in this subsection, we assume that $Y^{N}_{j,t}$ is a continuous random variable with a probability density function (PDF). Under this scenario, we suppose that the density of $Y^{N}_{0,t}$ can be replicated by a weighted sum of the densities of $Y^{N}_{j,t}$:
@@
We next discuss estimating SC weights that satisfy (3). One approach employs optimal transport to minimize the Wasserstein distance (Gunsilius, 2023). In this study, we propose a simpler estimation procedure, as the approach of Gunsilius (2023) involves more complexity than the original SCMs.
Because the mixture model (3) holds, we obtain:
$\mathbb{E}_{0,t}\left[\exp\left(\sqrt{-1}\xi Y^{N}_{0,t}\right)\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[\exp\left(\sqrt{-1}\xi Y^{N}_{j,t}\right)\right]\quad\forall t\in\mathcal{T},\ \forall\xi\in\mathbb{R},$
where $\sqrt{-1}$ denotes the imaginary unit. From this, we can estimate $\bm{w}^{*}$ by considering the upper bound of the difference between the expected exponential transforms of $Y^{N}_{0,t}$ and the corresponding weighted sum for $Y^{N}_{j,t}$. Specifically, we have
@@
$=\int_{\xi\in(-h,h)}\Biggl{|}\sum_{\gamma\in\{0,1,2,\dots,G\}}\frac{\sqrt{-1}^{\gamma}\xi^{\gamma}}{\gamma!}\left(\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right)+\epsilon\left|\xi^{G}\right|\Biggr{|}\mathrm{d}\xi$
@@
thus yields an accurate estimate of $\bm{w}^{*}$ that enforces
@@
This approach can be seen as a variant of the method of moments, where we match all moments of $Y^{N}_{0,t}$ with those of the weighted sum of the untreated units.
@@ 5.2 Core Assumptions
The outcomes $(Y^{N}_{j,t})_{j\in\mathcal{J},t\in\mathcal{T}}$ are bounded.
@@
Assumption 5.1 ensures that $Y^{N}_{j,t}$ has a finite $G$-th moment for any $G\in\mathbb{N}$. Assumption 5.2 has been employed in earlier work, including Wan et al. (2018), Gunsilius (2023), and Shi et al. (2022). Wan et al. (2018) notes that SCMs effectively rely on Assumption 5.2. Gunsilius (2023) similarly assumes (3) and estimates $\bm{w}^{*}$ by minimizing the Wasserstein distance, while Shi et al. (2022) introduces fine-grained models that, combined with Assumption 5.2, can justify standard SCMs. We provide further discussion of these fine-grained models in Section 8.1.
@@ 5.3 Estimation of SC Weights
Under the mixture model assumption, we propose estimating SC weights by matching characteristic functions (CHFs). We call our approach the Density Matching SCM (DMSCM).
Let $h\in(0,1)$ be a fixed parameter. We then estimate the SC weights as
$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\coloneqq\left(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0},G}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|.$ (4)
Using $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}$, we can then estimate both the treatment effect and the distributional treatment effect as follows.
5.3.1 Treatment Effect Estimation
Once $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}$ is obtained, we estimate the treatment effect similarly to standard SCMs:
$\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0},G}Y_{j,t}.$ (5)
We call this estimator the Density Matching SCM (DMSCM) estimator. Algorithm 1 summarizes our proposed estimator and also includes the D2MSCM estimator introduced later.
5.3.2 Distributional Treatment Effect Estimation
To estimate the distributional treatment effect, we can approximate the counterfactual distribution $p_{0,t}(y)$ via bootstrap. Specifically, we resample $L>1$ observations $\{Y^{l,\dagger}_{0}\}_{l=1}^{L}$ from $\{Y_{0,t},Y_{1,t},\dots,Y_{J,t}\}_{t\in\mathcal{T}_{1}}$ as follows: for each $l\in\mathcal{L}\coloneqq\{1,2,\dots,L\}$,
  1. 1.
  2. 2.
    Choose $Y^{l,\dagger}_{0}$ from $\left(Y^{l,\dagger}_{1},\dots,Y^{l,\dagger}_{J}\right)$ with probability $\left(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0}}\right)_{j\in\mathcal{J}^{U}}$.
Because our primary focus is often on $\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p_{0,t}(y)$, this bootstrap procedure enables us to approximate the distribution of the counterfactual outcomes.
Algorithm 1 DMSCM.
@@ 5.3.2 Distributional Treatment Effect Estimation
  Obtain trained weights $\widehat{\bm{w}}^{\mathrm{DMSCM}}$ as (4).
@@ 6 Convergence Analysis
Assume that the density function of $\varepsilon_{j,t}$ exists and is bounded for each $j\in\mathcal{J}$ and $t\in\mathcal{T}_{0}$, and that the stochastic process $\{\varepsilon_{j,t}\}_{t=1}^{T_{0}}$ is stationary, strongly mixing, with the sum of mixing coefficients bounded by $M$.
@@
Suppose that Assumptions 5.15.2 and 6.1 hold. Fix $h\in(0,1)$. Define
$\widetilde{\Phi}^{\dagger}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|.$
@@
$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}\quad\text{as }T_{0}\to\infty.$
@@
$\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}\xrightarrow{\mathrm{p}}\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)\quad\text{as }T_{0}\to\infty\text{ and }G\to\infty.$
The proof is shown in Appendix A. Note that this result also applies to $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ as a special case. Consequently, under the respective assumptions, $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ and $\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}$ are asymptotically unbiased.
7 Inference
This section presents methods for statistical inference using our proposed DMSCM estimator.
7.1 Conformal Inference on the Treatment Effect
We consider statistical inference for the treatment effect using our proposed DMSCM and D2SCM, focusing on testing a sharp null hypothesis, $H_{0}:\tau_{0,t}=\alpha_{t}$ for some $t\in\mathcal{T}_{t}$ and $\bm{\alpha}=(\alpha_{t})_{t\in\mathcal{T}_{1}}$. Specifically, we adopt the conformal inference approach of Chernozhukov et al. (2021), which has been used by, for example, Ben-Michael et al. (2021) and Ferman and Pinto (2021).

### v5

@@ 5.1 Identification and Estimation Strategy
Section 4 and the results in Ferman & Pinto (2021) indicate that combining
@@
Identification. We begin by positing that the distribution of the target unit’s outcome $Y^{N}_{0,t}$ can be expressed as a linear combination of the distributions of $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. For simplicity, in this subsection, we assume that $Y^{N}_{j,t}$ is a continuous random variable with a probability density function (PDF). Under this scenario, we suppose that the density of $Y^{N}_{0,t}$ can be replicated by a weighted sum of the densities of $Y^{N}_{j,t}$:
@@
Estimation. We next discuss how to estimate the SC weights that satisfy (3). One approach employs optimal transport to minimize the Wasserstein distance (Gunsilius, 2023). In this study, we propose a simpler estimation procedure, as the approach of Gunsilius (2023) is more complex than that of the original SCMs.
There are various ways to estimate $(w^{*}_{j})_{j\in\mathcal{J}^{U}}$. We consider the following:
  • Distribution divergence (distance) minimization.
  • Characteristic function matching.
  • Moment matching.
A straightforward approach to estimating $(w^{*}_{j})_{j\in\mathcal{J}^{U}}$ is to minimize the divergence or distance between two distributions whose densities are given by $p^{N}_{0,t}(y)$ and $\sum_{j\in\mathcal{J}^{U}}w_{j}p^{N}_{j,t}(y)$, given $(w_{j})_{j\in\mathcal{J}^{U}}$. For the divergence or distance measures, we may use, for example, the Kullback-Leibler (KL) divergence or the Wasserstein distance. The KL divergence can be estimated nonparametrically using density ratio techniques (assuming the densities exist) (Sugiyama et al., 2012; Kato & Teshima, 2021). Alternatively, we may estimate the weights by minimizing the maximum mean discrepancy.
The second approach is to match characteristic functions. Since the mixture model (3) holds, we have $\mathbb{E}_{0,t}\left[\exp\left(\sqrt{-1}\xi Y^{N}_{0,t}\right)\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[\exp\left(\sqrt{-1}\xi Y^{N}_{j,t}\right)\right]$ for all $t\in\mathcal{T}$ and $\xi\in\mathbb{R}$. where $\sqrt{-1}$ denotes the imaginary unit. The LHS is the characteristic function (CHF) of $p^{N}_{0,t}(y)$, and the RHS is that of $\sum_{j\in\mathcal{J}^{U}}w_{j}p^{N}_{j,t}(y)$, given $(w_{j})_{j\in\mathcal{J}^{U}}$. Since CHFs uniquely determine the corresponding distributions, we can estimate $(w^{*}_{j})_{j\in\mathcal{J}^{U}}$ by matching the CHFs. Such matching can be implemented using the method of moments, for instance, the approach proposed in Carrasco & Kotchoni (2017).
The last approach is to match the moments of the distributions $p^{N}_{0,t}(y)$ and $\sum_{j\in\mathcal{J}^{U}}w_{j}p^{N}_{j,t}(y)$. Intuitively, if we find $(w_{j})_{j\in\mathcal{J}^{U}}$ such that all the moments of the distributions match, we can interpret it as an estimate of $(w^{*}_{j})_{j\in\mathcal{J}^{U}}$. This idea can also be motivated from the perspective of CHF matching, as noted in Remark 5.2. Note that to justify this approach, we must ensure that the moments uniquely determine the distributions. This is the classical moment problem (Schmüdgen, 2020), which has been extensively studied. It is known that certain conditions are required to guarantee uniqueness.
Among the three approaches, we focus on the last one, moment matching. As we show later, it is sufficient to estimate $(w^{*}_{j})_{j\in\mathcal{J}^{U}}$ by matching sample means of polynomials of $Y_{a,j}$. Compared to the other approaches, the implementation of moment matching is significantly simpler. Note that it cannot be applied to all general distributions without additional assumptions since the moments may not be uniquely determined for each distribution. However, if we assume boundedness of $Y_{a,j}$, we can ensure that the distributions are uniquely determined by their moments (Hausdorff, 1921). In this study, for simplicity, we adopt this moment matching approach under the assumption of bounded random variables.
Remark 5.1 (Distribution SCMs using divergence or distance minimization.).
Since we are interested in finding $(w_{j})_{j\in\mathcal{J}^{U}}$ such that two distributions match, it is sufficient to use a measure that guarantees the distributions are identical when the divergence or distance is zero. That is, we do not need to consider symmetry or deficient support, which often motivate the use of the Wasserstein distance instead of the KL divergence. Therefore, while Gunsilius (2023) introduces distributional SCM using the Wasserstein distance, we believe that there is no strong advantage to using it over other divergence or distance measures.
Remark 5.2 (From CHF matching to moment matching).
We can estimate $\bm{w}^{*}$ by bounding the difference between the CHFs of $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{j,t}$. Specifically, we have
@@
$=\int_{\xi\in(-h,h)}\Biggl|\sum_{\gamma\in\{0,1,2,\dots,G\}}\frac{\sqrt{-1}^{\gamma}\xi^{\gamma}}{\gamma!}\left(\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right)+\epsilon|\xi^{G}|\Biggr|\mathrm{d}\xi$
@@
thus yields an accurate estimate of $\bm{w}^{*}$ that ensures
@@
This approach can be viewed as a variant of the method of moments, where we match all moments of $Y^{N}_{0,t}$ with those of the weighted sum of the untreated units.
Remark 5.3 (Reports in Gunsilius (2023)).
Distributional SCMs are also considered in Gunsilius (2023). However, the author does not emphasize the fact that distributional SCMs can estimate the treatment effect in an asymptotically unbiased manner, whereas original SCMs such as Abadie & Gardeazabal (2003) yield biased estimators. For example, Gunsilius (2023) states that “the distributional synthetic controls method finds optimal weights that replicate all moments of the target distribution as closely as possible. In contrast, applying the classical method to averages of the distributions will find the optimal weights based on replicating the first moment.” This claim may be accurate, as the classical method cannot yield unbiased and consistent estimates of the optimal weights for replicating the first moment unless, for example, we assume that the outcomes are nonrandom variables.
@@ 5.2 Core Assumptions
The outcomes $(Y^{N}_{j,t})_{j\in\mathcal{J},t\in\mathcal{T}}$ are bounded as $Y^{N}_{j,t}\in[0,1]$.
@@
Assumption 5.4 ensures that $Y^{N}_{j,t}$ has a finite $G$-th moment for any $G\in\mathbb{N}$. If $Y^{N}_{j,t}$ is bounded but not in $[0,1]$, we can normalize $Y^{N}_{j,t}$ to satisfy this assumption without loss of generality. Assumption 5.5 has been employed in earlier work, including Wan et al. (2018), Gunsilius (2023), and Shi et al. (2022). Wan et al. (2018) notes that SCMs effectively rely on Assumption 5.5. Gunsilius (2023) similarly assumes (3) and estimates $\bm{w}^{*}$ by minimizing the Wasserstein distance, while Shi et al. (2022) introduces fine-grained models that, combined with Assumption 5.5, can justify standard SCMs. We provide further discussion of these fine-grained models in Section 6.1.
@@ 5.3 Estimation of SC Weights
Under the mixture model assumption, we propose estimating SC weights by matching moment functions. We refer to our approach as the Momemnt Matching SCM (MMSCM).
Let $v_{\gamma}\in(0,\infty)$ for $\gamma=1,2,\dots,G$ be a fixed weight sequence. We then estimate the SC weights as
$\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}\coloneqq\left(\widehat{w}^{\mathrm{MMSCM}}_{j,T_{0},G}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}v_{\gamma}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|.$ (4)
We may use any weights $v_{\gamma}\in(0,\infty)$ for $\gamma=1,2,\dots,G$. For example, based on Remark 5.2, we can choose $v_{\gamma}=\frac{2h^{\gamma+1}}{(\gamma+1)!}$. In this case, we estimate $(w^{*}_{j})_{j\in\mathcal{J}^{U}}$ as
$\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}\coloneqq\left(\widehat{w}^{\mathrm{MMSCM}}_{j,T_{0},G}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|.$
Using $\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}$, we can then estimate both the treatment effect and the distributional treatment effect, as discussed in the following subsections.
5.4 Treatment Effect Estimation
Once $\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}$ is obtained, we estimate the treatment effect similarly to standard SCMs:
$\widehat{\tau}^{\mathrm{MMSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{MMSCM}}_{j,T_{0},G}Y_{j,t}.$ (5)
We call this estimator the Moment Matching SCM (MMSCM) estimator. Algorithm 1 summarizes our proposed estimator.
Remark 5.6 (Multiple treatments).
We can extend our method to the case with multiple treatments $K\geq 2$. Assume that there are units $\widetilde{J},\cdots,-1$ that receive treatments $\widetilde{J},\cdots,-1$, while unit 0 receives treatment $1$. In such cases, how we assume the DGP of the outcomes is crucial. For example, if we assume that the outcomes of the treated units follow the same distribution, it is sufficient to take the average over the treated group; that is, we solve $\min_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}v_{\gamma}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\frac{1}{\widetilde{J}+1}\sum_{\widetilde{j}\in\{0,-1,\dots,\widetilde{J}\}}\left(Y^{N}_{\widetilde{j},t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|$. We can also consider other situations in this setting, such as differential treatment timing. These extensions remain an open issue in the general context of SCM, and are not limited to the method proposed in this study.
Algorithm 1 MMSCM.
@@ 5.4 Treatment Effect Estimation
 Obtain trained weights $\widehat{\bm{w}}^{\mathrm{MMSCM}}$ as (4).
@@ 5.5 Convergence Analysis
The stochastic process $\{\varepsilon_{j,t}\}_{t=1}^{T_{0}}$ is stationary, strongly mixing, with the sum of mixing coefficients bounded by $M$.
@@
Suppose that Assumptions 5.45.5 and 5.7 hold. Let $v_{\gamma}\in(0,\infty)$ for $\gamma=1,2,\dots,G$ be a fixed weight sequence. Define
$\widetilde{\Phi}^{\dagger}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}v_{\gamma}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|.$
@@
$\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}\quad\text{as }T_{0}\to\infty.$
@@
$\widehat{\tau}^{\mathrm{MMSCM}}_{0,t}\xrightarrow{\mathrm{p}}\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)\quad\text{as }T_{0}\to\infty\text{ and }G\to\infty.$
The proof is shown in Appendix A. Note that this result also applies to $\widehat{\tau}^{\mathrm{MMSCM}}_{0,t}$ as a special case.
Consequently, because $\mathbb{E}\left[\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right]=0$, $\widehat{\tau}^{\mathrm{MMSCM}}_{0,t}$ is asymptotically unbiased.
Relationship to the classical SCMs. We next consider the relationship between our MMSCM and the classical SCMs proposed in Abadie & Gardeazabal (2003). Abadie & Gardeazabal (2003) assumes that $Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$, which corresponds to a specific case of our mixture model assumption, $F_{Y^{N}_{0,t}}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F_{Y^{N}_{j,t}}(y)\quad\forall y\in\mathbb{R}$. Therefore, under the assumptions in Abadie & Gardeazabal (2003), by using the MMSCMs, we can estimate the weights in an asymptotically unbiased way.
Corollary 5.9.
Suppose that Assumptions 5.4 and 5.7 hold. Assume that
$Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$
holds. Let $v_{\gamma}\in(0,\infty)$ for $\gamma=1,2,\dots,G$ be a fixed weight sequence. Define
$\widetilde{\Phi}^{\dagger}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}v_{\gamma}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|.$
Then, it holds that
$\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}\quad\text{as }T_{0}\to\infty.$
$\widehat{\tau}^{\mathrm{MMSCM}}_{0,t}\xrightarrow{\mathrm{p}}\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)\quad\text{as }T_{0}\to\infty\text{ and }G\to\infty.$
Remark 5.10 (Section 4.1 in Gunsilius (2023)).
The proposed method is an extension of the classical method in a rigorous sense: if we apply it to probability measures supported on one point, that is, Dirac measures of the form $\delta_{y}(A)$, taking the value $1$ if $y\in A$ and 0 otherwise, then we obtain the same results as the classical method. This means the proposed method reduces to the classical estimator when we are only given aggregate values and not distributions.
5.6 Inference
This section presents methods for statistical inference using our proposed MMSCM estimator. We consider statistical inference for the treatment effect using our proposed MMSCM and D2SCM, focusing on testing a sharp null hypothesis, $H_{0}:\tau_{0,t}=\alpha_{t}$ for some $t\in\mathcal{T}_{t}$ and $\bm{\alpha}=(\alpha_{t})_{t\in\mathcal{T}_{1}}$. Specifically, we adopt the conformal inference approach of Chernozhukov et al. (2021), which has been used by, for example, Ben-Michael et al. (2021) and Ferman & Pinto (2021). Note that under our linear factor model, $\tau_{0,t}=Y^{N}_{0,t}-Y^{I}_{0,t}=\alpha_{t}$ holds, that is, the ATE equals the individual treatment effect. Therefore, our null is sharp. This formulation is also adopted in other works, such as Chernozhukov et al. (2021).